• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name K. Kumar

    Log entry time 15:24:01 on July17,2005

    Entry number 148343


    keyword=shift summary

    8:30 Ryan completes harp scans
    9:20 Complete checkout for current scan
    9:25  Ran spot ++
    9:25  Ask beam off and put in single foil carbon target
    9:29  Carbon target in
    9:30  MCC called to say they need to reboot raster IOC
    10:05 MCC restores 5 micro amp beam
    10:10 quick recheck with spot++
    10:30 Decided to move to detector alignment before doing current calibration
    10:35 Helium target in place
    10:40 Bob begins detector alignment with spectrometer DAQ
    10:45 0.5 microamps on He target
    10:50 Start with Det1:1850V, Det3: 1700V for kHz rates
    11:16 Change Det1: 1900V, Det 3: 1800V for 30 kHz rate
    11:45 Complete change of momentum from 2.7076 to 2.6530 in both arms
    13:00 Start phase II high current commissioning of helium target w/o sweeper
    13:15 Start run number 3656 for current calibration
    13:35  Ramp up to 10 microamps
    13:45  Ramp up to 20 microamps
    13:50  Ramp up to 25 microamps
    14:10 End run 3656
    14:12  Ask for 2 microamps
    14:20  37A on sweeper
    14:23  74A on sweeper
    14:28  111A on sweeper
    14:30   10 microamps, 111A on sweeper
    14:36   25 microamps, 111A on sweeper
    14:50   35 microamps, 111A on sweeper
    15:15   Still waiting for 35 microamps to stabilize. Talked to Eric Abkamaier 
    and he agreed to set the trip level at 70 microamps at this raster (using a 
    linear extrapolation). This should be sufficient for us. However, if we need to 
    change the set points again, he is willing to authorize this remotely after 
    talking to MCC.